OpenStudy (adam.v271):

In Exam 1 the very last question - in the space of all 3x3 matrices we should consider the subspace of all rref's...I suppose one of the possible answers is that rref's do not form a subspace, since they are not closed under multiplication (that would break the 'leading-ones-rule')...

OpenStudy (anonymous):

The point is that the RREF matrices themselves do not themselves form a subspace, but that they *span* a subspace. You're absolutely right that, say (and I use \(2\times2\) matrices for brevity) \[\mathrm{B}=\left\{\begin{bmatrix}1&0\\0&1\end{bmatrix},\begin{bmatrix}1&1\\0&0\end{bmatrix},\begin{bmatrix}1&0\\0&0\end{bmatrix}\right\}\]is not a subspace because this set of matrices is not closed under multiplication. But the question isn't concerned with whether \(\mathrm{B}\) itself is a subspace. Instead, we're interested with what subspace of all \(2\times2\) matrices is spanned by \(\mathrm{B}\) (and hence what a basis for this subspace might be - it happens to be \(\mathrm{B}\)).